Four Seasons Hair Salon is the only women's hair stylist in Lake View. A business consultant has estimated its monthly cost functions as follows:

TC = 2000 + 100 Q + 0.05 Q  2
MC = 100 + 0.1 Q

The firm is faced with the following demand:

Qd = 1000 - 4 P

e. What price should it charge if it were to maximize its revenue?
f. At what level of output is the firm's average total cost minimized? Be sure to show your work.
Answer #1

liven Tc z 2000 + 100 Q + 0.05% me= loopt olla and demand Qd = 1000 - 49 Q = 1000-4p 4P = 1000 - Q P = 1000m 250 -0.250 4 manminimize Average tołal Cait d Ate) zo od ♡ 000 +100 +0.050) 20 dog 2000 toe05 ao 2ooo F 4000 = 70.05 2000 so.0502 Q² = 2050 0